Find a number which is 150 greater than its opposite.
Answer by funmath(2873)

Find a number which is 150 greater than its opposite.
Let a number be: x
Then it's opposite is: -x
x=-x+150
x+x=x-x+150
2x=150
2x/2=150/2
x=75
The number that is 150 greater than it's opposite is: 75
Happy Calculating!!!!